i) He owes $80,000 for his house which he believes is worth $150,000
ii) His car is worth $20,000 against which there there is $2,000 on the remaining bank loan.
iii) his stock protfolio has risen to $50,000
iv) he has a $10,000 bank balance in his bank account which is earning him 1.2% annual interest rate
v) and the value of this belonging is $45,000. he has just received his monthly paycheck for $6,000 and he is trying to decide about taking a vacation and whether or not to pay off his car loan his monthly expenses are $3,000 which includes the interest expense on his auto loan
vi)his two possible vacation choices are the bahamas for $2,000 or a local beach for $1,000
If he doesn't have enough money to cover all of his expenses for the month he will sell enough of his stock to cover the excess expenses.
a. use a spreadsheet to inpurt each of georges assets (i) to (v) in the first column, the value of these assets in the second column; and the liabilities (if any) against those assets in the third column. In the fourth column compute the net asset value of each of the assets. Total the fourth column to determine GEorge's NEt worth.
b. Compute the additional net income that George will have from his paycheck plus the interest rate interest on his bank account minus the monthly expenses Use this info to answer parts c through f below.
c. Reapeat part (a) for the end of the month assuming George does not take a vacation anys off his auto loans.
d. Repeat part (a) for the end of the month assuming George takes the Bahamas vacation and only pays & 1,000 on the principal of the auto loan.
e. Repeat part (a) for the end of the month assuming that George takes the local beach vacation and pays off his auto loan.
f. Repeat part (a) for the end of the month assuming George takes the Bahamas vacation and pays off the auto loan.
Answer #1

(a)

Assets Liabilities
Home: $150,000 Mortgage Payable: $80,000
Car: $20,000 Car Loan Payable: $2,000
Cash: $10,000
Belonging: $45,000
Unearned Interest: 0.012*10,000 = $120
Investments: $50,000

(b) Net income per month: $6,000 + $10 (interest) = $6,010.
(c) If George pays off his loan, income = $6,010 - $2,000 = $4,010.

Assets Liabilities
Home: $150,000 Mortgage Payable: $80,000
Car: $20,000
Cash: $10,000
Belonging: $45,000
Unearned Interest: 0.012*10,000 = $120
Investments: $50,000


(d) George takes the vacation and pays off $1,000. Income = $6,010 - $1,000 - $1,000 = $4,010.

Assets Liabilities
Home: $150,000 Mortgage Payable: $80,000
Car: $20,000 Car Loan Payable: $1,000
Cash: $10,000
Belonging: $45,000
Unearned Interest: 0.012*10,000 = $120
Investments: $50,000